Jeje is 8 years older than her brother Mon. The sum of their ages is 36. Find their present ages.
Answer by the_tutor2011(8)

J=M+8
M+J=36

M+J=36
M+M+8=36
2M+8=36
2M=28
M=14
J=M+8=14+8=22

Jeja = 14 yr
Mon = 22 yr